# Automating Legal Concept Interpretation with LLMs: Retrieval, Generation, and Evaluation

### v2

@@ Abstract
Legal articles often include vague concepts for adapting to the ever-changing society. Providing detailed interpretations of these concepts is a critical and challenging task even for legal practitioners. It requires meticulous and professional annotations and summarizations by legal experts, which are admittedly time-consuming and expensive to collect at scale. By emulating legal experts’ doctrinal method, we introduce a novel framework, ATRIE, using large language models (LLMs) to AuTomatically Retrieve concept-related information, Interpret legal concepts, and Evaluate generated interpretations, eliminating dependence on legal experts. ATRIE comprises a legal concept interpreter and a legal concept interpretation evaluator. The interpreter uses LLMs to retrieve relevant information from judicial precedents and interpret legal concepts. The evaluator uses performance changes on legal concept entailment, a downstream task we propose, as a proxy of interpretation quality. Automatic and multifaceted human evaluations indicate that the quality of our interpretations is comparable to those written by legal experts, with superior comprehensiveness and readability. Although there remains a slight gap in accuracy, it can already assist legal practitioners in improving the efficiency of concept interpretation.
@@
Automating Legal Concept Interpretation with LLMs:
@@
Kangcheng Luo1thanks: These authors contributed equally to this work., Quzhe Huang11footnotemark: 1, Cong Jiang211footnotemark: 1, Yansong Feng1thanks: Corresponding author. 1 Wangxuan Institute of Computer Technology, Peking University 2 Peking University luokangcheng@stu.pku.edu.cn    huangquzhe977@gmail.com {jiangcong,fengyansong}@pku.edu.cn
@@ 1 Introduction
Interpreting legal concepts is always essential since laws are often vague Endicott (2000) and open-textured Hart and Green (2012) to cover diverse real-world situations. For legal professionals, accurate interpretation is the foundation of fair judgments Barak (2005). For laypeople, it determines whether they can understand and comply with the law, guiding their daily lives and decisions Dworkin (1982). As shown in Figure 1, Theft in a dwelling is usually punished more severely than common theft. But what exactly is a “dwelling”? Is a school dormitory, tent, or motorhome a dwelling? Without clear interpretation, the law risks inconsistent application, undermining justice and public trust Smits (2017).
@@
However, interpreting legal concepts is far from easy. The legal system has invested great human effort and resources into doctrinal legal research Tiller and Cross (2006) to interpret the law. The doctrinal method of legal experts for writing legal concept interpretation begins with extensively reading a large volume of previous legal cases, books, papers, and other concept-related materials to find valuable information Yung-chin Su (2024). Then, they summarize past experience on detailed applications of these vague legal concepts. However, there are still several challenges: (1) Time-consuming: Legal professionals must browse countless texts and cases to build a reliable interpretation. Despite advances in legal research tools, this remains a labor-intensive task that is not fully automated VanGestel and Micklitz (2011). (2) Untimely: New cases continue to emerge at an increasing rate as society and technology progress. However, traditional methods rely on manual case-by-case reading to update interpretations, which is usually far behind judicial practice Van Hoecke (2011). (3) Incomplete and Subjective: Interpretations are limited by human capability. It is impossible to cover all existing cases, and interpretations remain incomplete. Moreover, when selecting cases from the overall case pool, humans may unconsciously or even intentionally introduce their own biases Farnsworth et al. (2011).
@@
Previous studies have attempted to use LLMs to interpret legal concepts to alleviate the burden on human experts. Savelka et al. (2023) utilize GPT-4 to interpret open-textured legal concepts from statutory articles based on expert-annotated valuable sentences from case law. However, this work fails to address the above challenges because of the dependence on legal experts to (1) annotate concept-related valuable sentences from extensive volumes of case law and (2) evaluate the quality of LLM-generated legal concept interpretations.
@@
Inspired by legal experts’ doctrinal method, we introduce ATRIE, an automatic framework for interpreting legal concepts and evaluating the generated interpretations without legal experts’ intensive involvement. ATRIE comprises a legal concept interpreter and a legal concept interpretation evaluator. The interpreter employs a Retrieval-Augmented Generation (RAG) framework Lewis et al. (2020); Guu et al. (2020). It leverages LLMs to retrieve comprehensive and concept-related information from a vast database of past cases, and then generates concept interpretations based on this information. The evaluator is based on our proposed downstream task, called Legal Concept Entailment (LCE), which assesses models’ understanding of legal concepts. We provide a specific LLM with different concept interpretations as references and test how its performance changes on the LCE task, using this as a proxy for the quality of concept interpretation. We recruit a legal expert to select 16 typical vague legal concepts and construct an LCE dataset to validate the effectiveness of our framework. Our contributions are as follows:
@@
  • We propose a novel automatic framework for legal concept interpretation, which mimics doctrinal legal methods used by legal experts and eliminates experts’ involvement.
@@
  • Automatic and human evaluations demonstrate that our generated concept interpretations not only help LLMs better understand vague concepts but also achieve high quality comparable to those written by legal experts.
@@ 2 Related Works
Legal interpretation has been a longstanding challenge in the field of legal NLP Nyarko and Sanga (2022). Initially, rule-based methods Waterman and Peterson (1981); Paquin et al. (1991) provide users with tribunal decisions and doctrinal works to establish the meaning of open-textured legal concepts in specific contexts. With the advancement of deep learning, research Šavelka and Ashley (2021a, b) uses pre-trained language models to retrieve sentences from legal cases that are useful to explain legal concepts.
@@
Previous works rely on legal experts to annotate concept-related information or evaluate generated interpretations. As a result, they fail to address the challenges mentioned earlier. Therefore, we introduce an automatic framework for retrieving concept-related information, interpreting legal concepts, and evaluating generated interpretations.
3 Preliminaries
In this work, we rely exclusively on previous legal cases as reference materials to interpret vague concepts in the articles. We use cases because they are the most concrete and fundamental sources; books and papers often cite cases to support their arguments. Formally, we define it as follows. Given a legal article $a$ and a vague concept $c$ within it, the task is to generate a legal interpretation $e$ for concept $c$, detailing the circumstances under which $c$ applies or not.
@@ 4 Legal Concept Interpreter
Following the method of legal experts, our legal concept interpreter summarizes the detailed applications of a given vague concept in judicial practice based on relevant case judgments. Specifically, it is composed of three parts (Figure 1): (1) Retrieve: Retrieve case judgments that mention the concept. (2) Filter&Extract: Select cases where the concept is analyzed in detail within the judgments and extract the reasons why the concept applies or not. (3) Interpret: Use LLMs to generate the interpretation of the concept based on the extracted reasons.
4.1 Retrieving case judgments
@@
To find case judgments helpful to interpret the vague concept, the first step is to retrieve those mention the concept. Formally, given a vague concept $c$ and the article $a$ to which $c$ belongs, we find all the case judgments citing the number of article $a$ from a case database. Then, we retrieve the cases that mention concept $c$ through exact string matching. All the retrieved cases form case set $\mathcal{D}_{0}$.
@@
Our case judgment database is constructed by collecting case judgments published on China Judgments Online [^1] https://wenshu.court.gov.cn/. It’s the largest public case judgment platform in China and the official website hosted by the Supreme People’s Court of China. Our database includes cases from 1985 to 2021, which ensures the source’s comprehensiveness.
@@
A case judgment typically contains five parts: Header, Facts, Court View, Verdict, and Conclusion  [^2] Details of the case judgment structure are in Appendix A.. Among them, the court view section explains the legal rationale and basis for the judgment. We use exact string matching to retrieve the case judgments that contain the vague concept in their court view. Legal terminology demands precision with fixed expressions that rarely permit alternative phrasings, so this approach ensures accuracy over fuzzy matching methods like dense retrieval.
4.2 Filtering relevant case judgments and extracting reasons
@@
In this step, we filter relevant cases—defined as those in which the court view sections provide detailed reasons why the vague concept applies to the case or not—and extract the reasons. Filtering relevant cases is essential, as some cases are relatively simple. Judges may not provide detailed discussions of the concept in the judgment, thus not contributing valuable insights for generating interpretations  [^3] We show an example of a judgment that mentions the concept only and a relevant case judgment that discusses the concept in detail in Appendix B..
@@
First, we use LLMs to filter the relevant cases from $\mathcal{D}_{0}$ [^4] All the prompts we use are shown in Appendix G.. Taking the court view as input, we require the LLM to determine whether it provides a detailed reason $r$ and extract this reason if provided. The reason $r$ should be a combination of original sentences from the court view. Next, we prompt LLMs to determine whether the concept applies to the case based on the court view, yielding a binary label $l$ (Yes/No). From this process, we obtain a refined case set $\mathcal{D}_{1}$ containing cases that discuss the concept in detail in the court view.
@@
Upon analyzing the labels within $\mathcal{D}_{1}$, we observe the proportion of positive cases (where $c$ applies to the case) far exceeds negative cases, with a ratio surpassing 10:1. This phenomenon could potentially be attributed to the exclusive inclusion of prosecuted and adjudicated cases in our sample. In judicial practice, only cases with substantial evidence supporting the prosecution are brought to court. As a result, the concept is more likely to apply to these cases, which leads to a higher proportion of positive examples. To comprehensively account for different situations when generating concept interpretations, we aim to ensure that both positive and negative examples receive adequate attention. Therefore, we only sample a subset of positive cases to construct a balanced dataset $\mathcal{D}$ and its corresponding reason set $\mathcal{R}$.
@@ 5 Legal Concept Interpretation Evaluator
Previous work has predominantly relied on human evaluation to evaluate the quality of the generated interpretations. We also conducted human evaluations, as detailed in Section 7. However, human evaluation is inherently subjective, and we aim to assess the quality of the generated concepts more objectively and quantitatively. Therefore, we design the legal concept interpretation evaluator based on a new task we propose, legal concept entailment. It enables an objective and reproducible comparison of different interpretations’ quality.
@@ 5.1 Legal Concept Entailment
If an interpretation of a concept is effective, it should help humans or models better determine whether the concept applies to previously unseen cases. Based on this assumption, we design the downstream task LCE. Given the fact description of a case relevant to the vague concept, the task is to determine whether the concept applies and provide a reason. We use a fixed LLM to perform this classification task. By incorporating different interpretations into the input, we can observe changes in the classification accuracy, which allows us to assess the quality of the interpretations. More accurate classification demonstrates higher-quality interpretation.
@@ 5.2 LCE Dataset
We recruit a legal expert with extensive judicial experience to identify 16 vague legal concepts in 14 legal articles (Appendix H). These concepts are typical and representative and frequently used in judicial practice. The statistical analysis reveals that, among all the cases in our database that cite these legal articles, 24.9% involve the corresponding vague concepts. Thus, we leverage them to demonstrate the effectiveness of our framework.
For each concept, we reuse the retrieval and filtering modules described in Sec 4.1 and 4.2 to collect relevant cases. These cases are challenging as the court views require detailed explanations of vague concepts. On average, 166 cases are selected for each concept, with a positive-to-negative case ratio 2:1. Detailed statistics are provided in Appendix H.
@@
The distinction between data annotation and LCE task lies in the input provided to the LLM. For annotation, the input is the court view, which contains explicit judgments made by judges and can be directly extracted as ground truth. In contrast, for the task itself, the input is the fact description, which lacks explicit judgments, requiring the LLM to perform reasoning to infer the entailment.
@@ 5.3 Evaluation Metrics
5.4 Method
This section introduces how our evaluator works. First, we generate the interpretations to be evaluated using our legal concept interpreter. To prevent data leakage, the cases used for generating interpretations do not overlap with the test dataset. Next, we prompt the LLM to perform the LCE task using the generated interpretations.
As shown in the right half of Figure 2, given a vague concept $c$ in a legal article $a$ and the fact description $f$ of a relevant case $d$, the LLM is prompted to analyze whether the concept $c$ applies to the fact $f$ based on the concept interpretation. Specifically, the LLM first generates a reason $\hat{r}$ and subsequently assigns a classification label $\hat{l}$.  [^5] Implementation details can be found in Appendix C.1.

### v3

# Automating Legal Interpretation with LLMs: Retrieval, Generation, and Evaluation

@@ Abstract
Interpreting the law is always essential for the law to adapt to the ever-changing society. It is a critical and challenging task even for legal practitioners, as it requires meticulous and professional annotations and summarizations by legal experts, which are admittedly time-consuming and expensive to collect at scale. To alleviate the burden on legal experts, we propose a method for automated legal interpretation. Specifically, by emulating doctrinal legal research, we introduce a novel framework, ATRIE, to address Legal Concept Interpretation, a typical task in legal interpretation. ATRIE utilizes large language models (LLMs) to AuTomatically Retrieve concept-related information, Interpret legal concepts, and Evaluate generated interpretations, eliminating dependence on legal experts. ATRIE comprises a legal concept interpreter and a legal concept interpretation evaluator. The interpreter uses LLMs to retrieve relevant information from previous cases and interpret legal concepts. The evaluator uses performance changes on Legal Concept Entailment, a downstream task we propose, as a proxy of interpretation quality. Automated and multifaceted human evaluations indicate that the quality of our interpretations is comparable to those written by legal experts, with superior comprehensiveness and readability. Although there remains a slight gap in accuracy, it can already assist legal practitioners in improving the efficiency of legal interpretation. [^1] The code and dataset are publicly available at https://github.com/lkc233/ATRIE
@@
Automating Legal Interpretation with LLMs:
@@
Kangcheng Luothanks: These authors contributed equally to this work., Quzhe Huang11footnotemark: 1, Cong Jiang11footnotemark: 1, Yansong Fengthanks: Corresponding author. Wangxuan Institute of Computer Technology, Peking University Institute for Artificial Intelligence, Peking University and Peking University Law School luokangcheng@stu.pku.edu.cn    huangquzhe977@gmail.com {jiangcong,fengyansong}@pku.edu.cn
@@ 1 Introduction
Interpreting the law is always essential since laws are often vague Endicott (2000) and open-textured Hart and Green (2012) to cover diverse real-world situations. For legal professionals, accurate interpretation is the foundation of fair judgments Barak (2005). For laypeople, it determines whether they can understand and comply with the law, guiding their daily lives and decisions Dworkin (1982). As shown in Figure 1, Theft in a dwelling is usually punished more severely than common theft. But what exactly is a “dwelling”? Is a school dormitory, tent, or motorhome a dwelling? Without clear interpretation, the law risks inconsistent application, undermining justice and public trust Smits (2017).
@@
However, interpreting the law is far from easy. The main method that legal systems have developed is doctrinal legal research, which aims to provide clear, systematic, and well-reasoned legal interpretations Tiller and Cross (2006). Doctrinal legal research involves legal experts extensively reading a large volume of previous legal cases, books, papers, and other related materials to find valuable information Yung-chin Su (2024). Then, they summarize experience on detailed applications of the law. However, there are still several challenges: (1) Time-consuming: Legal professionals must browse countless texts and cases to build a reliable interpretation. Despite advances in legal research tools, this remains a labor-intensive task that is not fully automated VanGestel and Micklitz (2011). (2) Untimely: New cases continue to emerge at an increasing rate as society and technology progress. However, traditional methods rely on manual case-by-case reading to update interpretations, which is usually far behind judicial practice Van Hoecke (2011). (3) Incomplete and Subjective: Interpretations are limited by human capability. It is impossible to cover all existing cases, and interpretations remain incomplete. Moreover, when selecting cases from the overall case pool, humans may unconsciously or even intentionally introduce their own biases Farnsworth et al. (2011).
@@
Previous studies have attempted to leverage LLMs for legal interpretation to mitigate the workload of human experts. Hoffman and Arbel (2024) and Engel and McAdams (2024) directly prompt LLMs for legal interpretation, asking LLMs to give the ordinary meaning of legal concepts. Savelka et al. (2023) utilize GPT-4 to interpret open-textured legal concepts from statutory articles based on expert-annotated valuable sentences from case law. These works have focused on interpreting legal concepts, as concepts represent the most important component of legal interpretation. Following prior research, we also select legal concept interpretation as the central focus of our work. However, previous works fail to address the above challenges because of the dependence on legal experts to (1) annotate concept-related valuable sentences from extensive volumes of case law and (2) evaluate the quality of LLM-generated legal concept interpretations.
@@
Inspired by doctrinal legal research, we introduce ATRIE to AuTomatically Retrieve concept-related information, Interpret legal concepts, and Evaluate generated interpretations without legal experts’ intensive involvement. ATRIE comprises a legal concept interpreter and a legal concept interpretation evaluator. The interpreter employs a Retrieval-Augmented Generation (RAG) framework Lewis et al. (2020); Guu et al. (2020). It leverages LLMs to retrieve comprehensive and concept-related information from a vast database of previous cases, and then generates concept interpretations based on this information. The evaluator is based on our proposed downstream task, called Legal Concept Entailment (LCE), which assesses models’ understanding of legal concepts. Provided with different concept interpretations as references, the performance of a specific LLM on the LCE task serves as a proxy for the quality of concept interpretation. We recruit a legal expert to select 16 typical vague legal concepts and construct an LCE dataset to validate the effectiveness of our framework. Our contributions are as follows:
@@
  • We propose a novel automated framework for legal concept interpretation, which mimics doctrinal legal methods used by legal experts and eliminates experts’ involvement.
@@
  • Automated and human evaluations demonstrate that our generated concept interpretations not only help LLMs better understand vague concepts but also achieve high quality comparable to those written by legal experts.
@@ 2 Related Works
2.1 Legal Interpretation
Legal interpretation is the process of identifying the meaning of legal texts Holmes (1898). It has been a longstanding challenge in the field of legal NLP Nyarko and Sanga (2022). Initially, rule-based methods Waterman and Peterson (1981); Paquin et al. (1991) provide users with tribunal decisions and doctrinal works to establish the meaning of open-textured legal concepts in specific contexts. With the advancement of deep learning, researchs Šavelka and Ashley (2021a, b) use pre-trained language models to retrieve sentences from legal cases which are useful to explain legal concepts.
@@ 2.1 Legal Interpretation
Previous approaches depend on legal experts to annotate concept-related information and evaluate generated interpretations. This reliance on manual expert input prevents them from addressing the aforementioned challenges. To overcome this, we introduce an automated framework that retrieves concept-related information, interprets legal concepts, and evaluates the resulting interpretations.
2.2 Doctrinal Legal Research
Doctrinal legal research, which dates back to the Roman Empire Van Hoecke and Ost (1998), is a traditional and systematic method of analyzing legal texts, case law, and principles Pound (1913). The goal is to identify, explain, and systematize legal rules, and sometimes to predict how courts might apply them in the future Tiller and Cross (2006). This method is widely accepted in courts and legal scholarship as the standard way to work with legal texts around the world Bhat (2019). Many legal tasks, including legal interpretation, mainly depend on doctrinal legal research Majeed et al. (2023). Doctrinal legal research provides a clear approach to analyze the meaning of unclear or vague legal texts. Traditionally, this work is done by human experts. It often relies on significant human effort to review extensive cases, identify patterns, and extract insights. This process is time-consuming and labor-intensive Hutchinson and Duncan (2012). Yung-chin Su (2024) suggests that the task within doctrinal legal research of reading cases and extracting theoretical insights could potentially be performed by Legal AI. Inspired by the process of doctrinal legal research, our work explores how LLMs can automate legal interpretation, making it faster and more scalable.
3 Task Definition
In this work, we rely exclusively on previous cases as reference materials to interpret vague concepts in the articles. We use cases because they are the most concrete and fundamental sources. Books and papers often cite cases to support their arguments.
We formally define the task as follows: Given a legal article $a$ and a vague concept $c$ contained within, the objective is to generate a legal interpretation $e$ for $c$ based on previous cases, clarifying the conditions of its applicability. For example, as shown in Figure 1, our goal is to address the vague concept of "Dwelling" within Article 264 of the Criminal Law. We aim to leverage previous cases to interpret the circumstances under which "Dwelling" applies.
@@ 4 Legal Concept Interpreter
Following the method of legal experts, our legal concept interpreter summarizes the detailed applications of a given vague concept in judicial practice based on relevant cases. Specifically, it is composed of three parts (Figure 1): (1) Retrieve: Retrieve cases that mention the concept. (2) Filter&Extract: Select cases where the concept is analyzed in detail within the cases and extract the reasons why the concept applies or not. (3) Interpret: Use LLMs to generate the interpretation of the concept based on the extracted reasons.
4.1 Retrieving Cases
@@
To find cases helpful to interpret the vague concept, the first step is to retrieve those mention the concept. Formally, given a vague concept $c$ and the article $a$ to which $c$ belongs, we find all the cases citing the number of article $a$ from a case database. Then, we retrieve the cases that mention concept $c$ through exact string matching. All the retrieved cases form case set $\mathcal{D}_{0}$.
@@
Our case database is constructed by collecting judgment documents published on China Judgments Online [^2] https://wenshu.court.gov.cn/. It’s the largest public case platform in China and the official website hosted by the Supreme People’s Court of China. Our database includes cases from 1985 to 2021, which ensures the source’s comprehensiveness.
@@
To further refine this retrieval process, we focus specifically on the most relevant section of the case document. A case typically contains five parts: Header, Facts, Court View, Verdict, and Conclusion  [^3] Details of the case structure are in Appendix A.. Among them, the court view section explains the legal rationale and basis for the judgment. We use exact string matching to retrieve the cases that contain the vague concept in their court views. Legal terminology demands precision with fixed expressions that rarely permit alternative phrasings, so this approach ensures accuracy over fuzzy matching methods like dense retrieval.
4.2 Filtering Relevant Cases and Extracting Reasons
@@
In this step, we filter relevant cases—defined as those in which the court view sections provide detailed reasons why the vague concept applies to the case or not—and extract the reasons. This filtering is necessary because cases lacking such detailed discussion offer few valuable insights for generating interpretations.  [^4] We show an example of a case that mentions the concept only and a relevant case that discusses the concept in detail in Appendix B.
@@
First, we use LLMs to filter the relevant cases from $\mathcal{D}_{0}$. [^5] All the prompts we use are shown in Appendix G. Taking the court view as input, we require the LLM to determine whether it provides a detailed reason $r$ and extract this reason if provided. The reason $r$ should be a combination of original sentences from the court view. Next, we prompt LLMs to determine whether the concept applies to the case based on the court view, yielding a binary label $l$ (Yes/No). From this process, we obtain a refined case set $\mathcal{D}_{1}$ containing cases that discuss the concept in detail in the court view.
@@
Upon analyzing the labels within $\mathcal{D}_{1}$, we observe the proportion of positive cases (where $c$ applies to the case) far exceeds negative cases, with a ratio surpassing 10:1. This phenomenon could potentially be attributed to the exclusive inclusion of prosecuted and adjudicated cases in China Judgments Online. In judicial practice, only cases with substantial evidence supporting the prosecution are brought to court. As a result, the concept is more likely to apply to these cases, which leads to a higher proportion of positive examples. To comprehensively account for different situations when generating concept interpretations, we aim to ensure that both positive and negative examples receive adequate attention. Therefore, we only sample a subset of positive cases to construct a balanced dataset $\mathcal{D}$ and its corresponding reason set $\mathcal{R}$.
@@ 5 Legal Concept Interpretation Evaluator
Previous work Savelka et al. (2023) has predominantly relied on human evaluation to evaluate the quality of the generated interpretations. We also conducted human evaluations, as detailed in Section 7. However, human evaluation is inherently subjective, and we aim to assess the quality of the generated concepts more objectively and quantitatively. Therefore, we design the legal concept interpretation evaluator based on a new task we propose, Legal Concept Entailment. It enables an objective and reproducible comparison of different interpretations’ quality.
@@ 5.1 Legal Concept Entailment
If an interpretation of a concept is effective, it should help humans or models better determine whether the concept applies to unseen cases. Based on this assumption, we design the downstream task Legal Concept Entailment (LCE). Given the fact description of a case relevant to the vague concept, the task is to determine whether the concept applies and provide a reason. We use a fixed LLM to perform this classification task. By incorporating different interpretations into the input, we can observe changes in the classification accuracy, which allows us to assess the quality of the interpretations. More accurate classification demonstrates higher-quality interpretation.
@@ 5.2 LCE Dataset
We recruit a legal expert with extensive judicial experience to identify 16 vague legal concepts in 14 legal articles (Appendix H). These concepts are representative and frequently used in judicial practice. The statistical analysis reveals that, among all the cases in our database that cite these legal articles, 24.9% involve the corresponding vague concepts. Thus, we leverage them to demonstrate the effectiveness of our framework.
For each concept, we reuse the retrieval and filtering modules described in Sec 4.1 and 4.2 to collect relevant cases. These cases are considered challenging because the court view provides detailed explanations of the vague concepts. On average, 166 cases are selected for each concept, with a positive-to-negative case ratio 2:1. Detailed statistics are provided in Appendix H.
@@
The distinction between data annotation and LCE task lies in the input provided to the LLM. For annotation, the input is the court view, which contains explicit judgments made by judges and can be directly extracted as ground truth. In contrast, for the LCE task itself, the input is the fact description, which lacks explicit judgments, requiring the LLM to perform reasoning to infer the entailment.
@@ 5.3 Evaluation Metrics
5.4 Evaluation Process
This section describes how our evaluator assesses the interpretations generated by our framework.
First, we generate the interpretations to be evaluated using our legal concept interpreter. To prevent data leakage, the cases used for generating interpretations do not overlap with the test dataset. Next, we prompt the LLM to perform the LCE task using the generated interpretations.
As shown in the right half of Figure 2, given a vague concept $c$ in a legal article $a$ and the fact description $f$ of a relevant case $d$, the LLM is prompted to analyze whether the concept $c$ applies to the fact $f$ based on the concept interpretation. Specifically, the LLM first generates a reason $\hat{r}$ and subsequently assigns a classification label $\hat{l}$.